# Oaken: Fast and Efficient LLM Serving with Online-Offline Hybrid KV Cache Quantization

### v1

@@ Abstract.
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXconference: The 52nd IEEE/ACM International Symposium on Computer Architecture; June 21–25, 2025; Tokyo, Japanisbn: 978-1-4503-XXXX-X/18/06submissionid: 123-A56-BU3
@@ 1. Introduction
Figure 1. Existing solutions for LLM inference serving systems plotted on the bandwidth-capacity trade-off space. The “effective bandwidth” and “effective capacity” represent the scale of data that can be transmitted to/from and stored on memory, respectively. We also delineate their corresponding throughputs (i.e., tokens/sec) using the colors presented on a 1D heatmap on the right side.
@@
As LLM inferencing is for multi-tenant environments, the serving systems batch multiple requests to parallelize the inference computation (Yu et al., 2022; Kwon et al., 2023; Sheng et al., 2023; Cheng et al., 2023a; Li et al., 2023; Dao et al., 2022; NVIDIA, 2023). While batching promises a significant throughput boost for operations where operand matrices can be shared across requests on GPUs and other AI accelerators, attention layers in transformer-based LLMs consist of un-batchable operations with request-specific, un-sharable operands, lowering hardware utilization. As the un-batchable operations cannot exploit on-chip data reuse, they produce enormous memory read, causing memory bandwidth to become the key system bottleneck.
@@
Prioritizing bandwidth over capacity: Currently, using HBM-equipped GPUs is the de-facto standard solution for LLM inference processing (Jouppi et al., 2023; Medina and Dagan, 2020; NVIDIA, 2020). While this approach achieves massive bandwidth, it often compromises capacity, with insufficient capacity forcing systems to scale out, which not only reduces core utilization but also increases the cost of building the system.
@@
Exploring LLM quantization strategies: One fundamental strategy for jointly addressing the conflicting objectives is to minimize the memory footprint required for LLM inferencing. To achieve this goal, a large body of prior work (Kim et al., 2023b; Hooper et al., 2024; Lee et al., 2024a; Lin et al., 2023; Kim et al., 2023a; Frantar et al., 2022b; Zadeh et al., 2020; Chee et al., 2024; Yao et al., 2022; Lin* et al., 2024; Zhao et al., 2024a; Liu et al., 2024; Yuan et al., 2023) have recently developed LLM-targeted quantization techniques. While these techniques successfully achieve significant reductions in bitwidth, they often prioritize minimizing bitwidth over effectively translating these reductions into practical inferencing speedups.
@@ 2.1. A Primer on LLM Inference
2.2. Batching for High-Throughput LLM Inferencing
@@ 2.3. LLM Quantization
Recently, as demands for LLM inference ever-increasing memory capacity, the community has explored solutions to mitigate this demand through quantization techniques that reduce the bit precision of weights or activations. Many of these studies focus on easing the memory pressure caused by model weights, which is particularly effective for accelerating small-batch inference (Kim et al., 2023b; Lin et al., 2023; Shao et al., 2023; Frantar et al., 2022a; Wei et al., 2022; Yuan et al., 2023; Wei et al., 2023; Liu et al., 2023). However, these methods achieve limited speedup for larger batches and long sequence lengths due to the KV cache, whose size scales with the batch size and sequence length. Approaches to quantize both weights and activations also face similar limitations (Xiao et al., 2023; Lee et al., 2023; Sheng et al., 2023; Dettmers et al., 2022). To this end, researchers have recently started directing their attention to KV cache quantization. Some of these works propose quantizing the keys and values of the attention layers on a per-vector basis (Hooper et al., 2024; Lin* et al., 2024; Liu et al., 2024), while others propose quantizing the KV cache into multiple vector groups of similar magnitudes, applying channel reordering technique (Zhao et al., 2024a; Lee et al., 2024b; Lin et al., 2024; Ashkboos et al., 2024).
@@ 3.1. Characterization of Batched LLM Inference
Single-instance inferencing. We analyze the bottlenecks of batched inference by examining the core utilization of the accelerator for both single and batched request scenarios. Figure 3(a) highlights the difference in core utilization between the prefill and generation phases when processing a single request. During the prefill phase, multiple cores are utilized to process multiple input tokens in a parallelized manner. However, the generation phase processes only one token from the previous iteration, making it highly sequential process that limits parallelization. Therefore, the accelerator can utilize only a few cores, resulting in high latency in generating all output tokens.
@@
Figure 6. Distribution of KV values: (a) Range of KV values from each decoder layer of OPT-6.7B and Llama2-7B models using Wikitext2 dataset; (b) Range of KV values from each decoder layer of Llama2-7B model using Wikitext2, PIQA, and Hellaswag dataset; (c) Distribution of the top 4% keys from the 6th decoder layer of Llama2-7B and OPT-6.7B models using Wikitext2 dataset.
@@ 3.3. KV Cache Quantization
Algorithmic approach to the memory wall. To tackle the memory wall challenge, the community has explored algorithmic approaches to reduce the demands on both conflicting resources fundamentally. Quantization is one such direction, widely recognized for its ability to alleviate both capacity and bandwidth bottlenecks simultaneously. As discussed in Section 2.3, many efforts focus on weight quantization and reducing the computation workload to accelerate LLM inference (Lee et al., 2024a; Dettmers et al., 2022; Frantar et al., 2022a; Park et al., 2023; Yuan et al., 2023; Xiao et al., 2023). However, our experimental results show that weight-only quantization has a limited impact on addressing the memory wall in the batched scenarios.
@@ 4.3. Threshold-based Online-Offline Hybrid Quantization
Oaken performs approximately a hundred offline inferences with sample input prompts to gather distribution information from the KV cache of each decoder layer. The four group thresholds are extracted during the profiling process from the key and value tensors of each inference run using topK operations, and theirr averages are computed for each decoder layer. These statistics are then used to establish group thresholds for the KV cache. Since different LLMs exhibit unique KV distributions, Oaken’s offline profiling should be performed separately for each model. However, as discussed in Section 4.1, KV cache distribution and the profiling process are independent of both the profiling dataset and future inputs. It effectively minimizes accuracy loss with only a small number of inferences, ensuring minimal and one-time overhead.
@@ 5.2. Oaken Compute Core
First, the 1 decomposer module partitions incoming activations into three quantization groups based on outlier thresholds determined offline. It then performs group shift for the outer and middle groups by subtracting their streamed thresholds from the values of these groups. Finally, the middle group is directed to the inlier quantization path, while outliers, whether from inner or outer groups, are routed to the outlier quantization path, with zeros inserted in the alternate path accordingly.
Both the 2 inlier and 3 outlier quantizer modules handle quantization for each key and value vector. Quantization scaling factors are dynamically computed during runtime, based on per-token min and max values for each group. After calculating the scaling factors, these two modules finally perform 4/5-bit uniform quantization. The quantized inlier and outlier values are merged using an OR gate and sent to the quantized dense matrix. The outlier module generates an index and a group flag for COO transformation. It employs a zero-remove shifter (Rhu et al., 2018; Wang et al., 2021) to implement fused dense-and-sparse encoding, optimizing memory usage.
@@
The 4 inlier dequantizer module buffers incoming dense data to synchronize with sparse data processed by the outlier dequantizer module. The 5 outlier dequantizer module handles sparse COO data by performing a zero-insert (Rhu et al., 2018; Wang et al., 2021) operation to restore the original data alignment. It identifies the original positions of fused outliers using the index and group information of the sparse data and inserts the necessary zeros accordingly. Both dequantizer modules then restore the data, which is buffered to be aligned with the outputs of the counterpart module. Finally, the outputs from both dequantizer modules are merged via an OR gate and forwarded to the processing unit.
@@ 5.3. Token-level Batch Scheduling
6. Evaluation
@@ 6.2. Experimental Results
Figure 12. (a) Accuracy and effective bits with varying quantization group ratios. (b) Throughput results on Llama2-13B model with a batch size of 16 when increasing total sequence length from 1K to 32K. The ratio of input and output length is set to 1:1.
@@
Figure 13. Latency breakdown for non-attention, attention, quantization and dequantization operations using Llama2-7B model across varying batch size.
@@
7. Related Work
@@
Acceleration for quantized model inference. There have been prior works on accelerating quantized neural network inference (Qin et al., 2023; Zeng et al., 2024; Reggiani et al., 2023; Chang et al., 2021; Song et al., 2020; Inci et al., 2023; Keller et al., 2022; Yang et al., 2022; Li et al., 2024a; Shen et al., 2024). Mokey (Zadeh et al., 2022) and Olive (Guo et al., 2023) use outlier-aware quantization methods for transformer-based LLMs. Tender (Lee et al., 2024b) quantizes KV cache as well as weights and activations, but its accuracy loss is significant. Oaken is an LLM inference acceleration technique optimized for KV cache quantization, offering a scalable solution while minimizing accuracy loss.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 52nd Annual International Symposium on Computer Architecture; June 21–25, 2025; Tokyo, Japanbooktitle: Proceedings of the 52nd Annual International Symposium on Computer Architecture (ISCA ’25), June 21–25, 2025, Tokyo, Japandoi: 10.1145/3695053.3731019isbn: 979-8-4007-1261-6/2025/06ccs: Computer systems organization Neural networksccs: Computing methodologies Neural networksccs: Computer systems organization Multicore architectures
@@ 1. Introduction
Figure 1. Existing solutions for LLM inference serving systems plotted on the bandwidth-capacity trade-off space. The “effective bandwidth” and “effective capacity” represent the scale of data that can be transmitted to/from and stored on memory, respectively. We also delineate their corresponding throughput (i.e., tokens/sec) using the colors presented on a 1D heatmap on the right side.
[…]
Existing solutions for LLM inference serving systems plotted on the bandwidth-capacity trade-off space. The “effective bandwidth” and “effective capacity” represent the scale of data that can be transmitted to/from and stored on memory, respectively. We also delineate their corresponding throughput (i.e., tokens/sec) using the colors presented on a 1D heatmap on the right side.
@@
As LLM inferencing is for multi-tenant environments, the serving systems batch multiple requests to parallelize the inference computation (Yu et al., 2022; Kwon et al., 2023; Sheng et al., 2023; Cheng et al., 2023a; Li et al., 2023; Dao et al., 2022; NVIDIA, 2023). While batching promises a significant throughput boost for operations where operand matrices can be shared across requests on GPUs and other AI accelerators, attention layers in transformer-based LLMs consist of un-batchable operations with request-specific, un-shareable operands, lowering hardware utilization. As the un-batchable operations cannot exploit on-chip data reuse, they produce enormous memory read, causing memory bandwidth to become the key system bottleneck.
@@
Prioritizing bandwidth over capacity: Currently, using HBM-equipped GPUs is the de-facto standard solution for LLM inference processing (Jouppi et al., 2023; Medina and Dagan, 2020; NVIDIA, 2020). While this approach achieves massive bandwidth, it often compromises capacity, forcing systems to scale out, which not only reduces core utilization but also increases the cost of building the system.
@@
Exploring LLM quantization strategies: One fundamental strategy for jointly addressing the conflicting objectives is to minimize the memory footprint required for LLM inferencing. To achieve this goal, a large body of prior work (Kim et al., 2023b; Hooper et al., 2024; Lee et al., 2024a; Lin et al., 2023; Kim et al., 2023a; Frantar et al., 2022; Zadeh et al., 2020; Chee et al., 2024; Yao et al., 2022; Lin* et al., 2024; Zhao et al., 2024a; Liu et al., 2024; Yuan et al., 2023) have recently developed LLM-targeted quantization techniques. While these techniques successfully achieve significant reductions in bitwidth, they often prioritize minimizing bitwidth over effectively translating these reductions into practical inferencing speedups.
@@
[…]
(a) Structure of LLM inference and decoder layer during the prefill and generation phases. (b) Operations in the multi-head attention layer, including activation-weight and activation-activation operations, during the generation phase of batched inference for three requests.
@@ 2.1. A Primer on LLM Inference
2.2. Batching for High-Throughput LLM Inference
@@ 2.3. LLM Quantization
Recently, as demands for LLM inference ever-increasing memory capacity, the community has explored solutions to mitigate this demand through quantization techniques that reduce the bit precision of weights or activations. Many of these studies focus on easing the memory pressure caused by model weights, which is particularly effective for accelerating small-batch inference (Kim et al., 2023b; Lin et al., 2023; Shao et al., 2023; Frantar et al., 2022; Wei et al., 2022; Yuan et al., 2023; Wei et al., 2023; Liu et al., 2023). However, these methods achieve limited speedup for larger batches and long sequence lengths due to the KV cache, whose size scales with the batch size and sequence length. Approaches to quantize both weights and activations also face similar limitations (Xiao et al., 2023; Lee et al., 2023; Sheng et al., 2023; Dettmers et al., 2022). To this end, researchers have recently started directing their attention to KV cache quantization. Some of these works propose quantizing the keys and values of the attention layers on a per-vector basis (Hooper et al., 2024; Lin* et al., 2024; Liu et al., 2024), while others propose quantizing the KV cache into multiple vector groups of similar magnitudes, applying channel reordering technique (Zhao et al., 2024a; Lee et al., 2024b; Lin et al., 2024; Ashkboos et al., 2024).
@@ 3. Motivation
[…]
Characteristic analysis of LLM inference for (a) single request and (b) batched multiple requests. (c) Utilization measurement during the generation phase with batched multiple requests using NVIDIA A100 GPU.
@@ 3.1. Characterization of Batched LLM Inference
Single-instance inferencing. We analyze the bottlenecks of batched inference by examining accelerator core utilization in both single and batched request scenarios. Figure 3(a) shows the difference in core utilization between the prefill and generation phases when processing a single request. During the prefill phase, multiple cores are utilized to process multiple input tokens in a parallelized manner. However, the generation phase processes only one token from the previous iteration, making it highly sequential process that limits parallelization. Therefore, the accelerator can utilize only a few cores, resulting in high latency in generating all output tokens.
@@
[…]
Throughput of accelerators equipped with HBM and LPDDR memory when using (a) Llama2-13B and (b) OPT-30B (OOM refers to “Out-of-Memory.”). (c) Accelerator specification with HBM and LPDDR memory.
@@ 3.2. Trade-off between Bandwidth and Capacity
[…]
(a) Memory usage breakdown to KV cache and model parameters of Llama2-13B model as batch size sweeps from 1 to 256. (b) Throughput comparison among no quantization, weight and KV cache quantization of Llama2-13B model inference. Experiment is conducted with LPDDR-NPU.
@@ 3.3. KV Cache Quantization
Algorithmic approach to the memory wall. To tackle the memory wall challenge, the community has explored algorithmic approaches to reduce the demands on both conflicting resources fundamentally. Quantization is one such direction, widely recognized for its ability to alleviate both capacity and bandwidth bottlenecks simultaneously. As discussed in Section 2.3, many efforts focus on weight quantization and reducing the computation workload to accelerate LLM inference (Lee et al., 2024a; Dettmers et al., 2022; Frantar et al., 2022; Park et al., 2024c; Yuan et al., 2023; Xiao et al., 2023). However, our experimental results show that weight-only quantization has a limited impact on addressing the memory wall in the batched scenarios.
@@ 4. Oaken’s KV Quantization Algorithm
Figure 6. (a) Range of KV values from each decoder layer of OPT-6.7B and Llama2-7B models using Wikitext2 dataset. (b) Range of KV values from each decoder layer of Llama2-7B model using Wikitext2, PIQA, and Hellaswag dataset. (c) Distribution of the top 4% keys from the 6th decoder layer of Llama2-7B and OPT-6.7B models using Wikitext2 dataset.
[…]
(a) Range of KV values from each decoder layer of OPT-6.7B and Llama2-7B models using Wikitext2 dataset. (b) Range of KV values from each decoder layer of Llama2-7B model using Wikitext2, PIQA, and Hellaswag dataset. (c) Distribution of the top 4% keys from the 6th decoder layer of Llama2-7B and OPT-6.7B models using Wikitext2 dataset.
@@ 4.1. Observations in KV Distribution
[…]
Oaken’s quantization algorithm consisting of three components: (a) threshold-based online-offline hybrid quantization, (b) group-shift quantization, and (c) fused dense-and-sparse encoding.
@@ 4.3. Threshold-based Online-Offline Hybrid Quantization
Oaken performs approximately a hundred offline inferences with sample input prompts to gather distribution information from the KV cache of each decoder layer. The four group thresholds are extracted during the profiling process from the KV cache of each inference run using topK operations, and their averages are computed for each decoder layer. These statistics are then used to establish group thresholds for the KV cache. As discussed in Section 4.1, Oaken’s offline profiling should be performed separately for each model, whereas the KV cache distribution and the profiling are independent of both the profiling dataset and future inputs. It effectively minimizes accuracy loss with only a small number of inferences, ensuring minimal and one-time overhead.
@@ 5.1. Architecture Overview
Figure 8. Overall Oaken accelerator architecture.
[…]
Overall Oaken accelerator architecture.
@@ 5.2. Oaken Compute Core
[…]
Quantization and dequantization engines of Oaken compute core.
@@
First, the decomposer module partitions incoming activations into three quantization groups based on outlier thresholds determined offline. It then performs group shift for the outer and middle groups by subtracting their streamed thresholds from the values of these groups. Finally, the middle group is directed to the inlier quantization path, while outliers, whether from inner or outer groups, are routed to the outlier quantization path, with zeros inserted in the alternate path accordingly.
Both the inlier and outlier quantizer modules handle quantization for each key and value vector. Quantization scaling factors are dynamically computed during runtime, based on per-token min and max values for each group. After calculating the scaling factors, these two modules finally perform 4/5-bit uniform quantization. The quantized inlier and outlier values are merged using an OR gate and sent to the quantized dense matrix. The outlier module generates an index and a group flag for COO transformation. It employs a zero-remove shifter (Rhu et al., 2018; Wang et al., 2021) to implement fused dense-and-sparse encoding, optimizing memory usage.
@@
The inlier dequantizer module buffers incoming dense data to synchronize with sparse data processed by the outlier dequantizer module. The outlier dequantizer module handles sparse COO data by performing a zero-insert (Rhu et al., 2018; Wang et al., 2021) operation to restore the original data alignment. It identifies the original positions of fused outliers using the index and group information of the sparse data and inserts the necessary zeros accordingly. Both dequantizer modules then restore the data, which is buffered to be aligned with the outputs of the counterpart module. Finally, the outputs from both dequantizer modules are merged via an OR gate and forwarded to the processing unit.
@@
[…]
Operations of memory management unit (MMU) for handling dense and sparse data.
@@ 5.3. Token-level Batch Scheduling
6. Evaluation
[…]
Throughput results of GPU baselines, LPU, Tender, and Oaken equipped with LPDDR and HBM across six LLMs. We sweep the batch size from 16 to 256. The input and output sequence lengths are set to 1K:1K.
@@ 6. Evaluation
[…]
Hardware specification of NVIDIA A100 GPU and Oaken equipped with either HBM or LPDDR memory. [Uncaptioned image]
@@ 6.1. Methodology
[…]
Perplexity results on Wikitext2, and zero-shot accuracy results on PIQA, Winogrande, and Hellaswag datasets with effective bitwidth of each quantization technique. [Uncaptioned image]
@@ 6.2. Experimental Results
Figure 12. (a) Accuracy and effective bits with varying quantization group ratios. (b) Latency breakdown for non-attention, attention, quantization and dequantization operations using Llama2-7B model across varying batch size.
[…]
(a) Accuracy and effective bits with varying quantization group ratios. (b) Latency breakdown for non-attention, attention, quantization and dequantization operations using Llama2-7B model across varying batch size.
@@
[…]
Accuracy and effective bits using the Llama2-7B model with varying number of groups and group ratios while keeping the total inner and outer group ratio at 10%. [Uncaptioned image]
Figure 13. Throughput results on Llama2-13B model with a batch size of 16 when increasing total sequence length from 1K to 32K. The ratio of input and output length is set to 1:1.
[…]
Throughput results on Llama2-13B model with a batch size of 16 when increasing total sequence length from 1K to 32K. The ratio of input and output length is set to 1:1.
@@
[…]
Real-world benchmark results for generation throughput on Llama2-13B and Mixtral-8x7B models evaluated with two LLM inference traces: Conversation and BurstGPT. We sweep the batch size from 16 to 128.
@@
7. Related Work
LLM quantization. Most prior work on LLM quantization focuses on weight and/or activation quantization (Lin et al., 2023; Shao et al., 2023; Frantar et al., 2022; Wei et al., 2022, 2023; Liu et al., 2023) to reduce inference computation costs. RPTQ (Yuan et al., 2023), SpinQuant (Liu et al., 2025), and QuaRot (Ashkboos et al., 2024) introduce transformation matrices for weight and activation quantization. SmoothQuant (Xiao et al., 2023) mitigates the quantization difficulty by transferring activation outlier scales to weights. SqueezeLLM (Kim et al., 2023b) applies dense-and-sparse quantization for storing weight outliers in full precision. However, many existing approaches overlook the KV cache, whose size scales with sequence length and batch size, often becoming a major bottleneck for latency and throughput in batched LLM inference. Oaken overcomes this issue by employing an offline-online KV cache quantization algorithm with a customized hardware module, achieving high throughput with minimal accuracy degradation.
@@ 7. Related Work
[…]
Area overhead analysis of compression and decompression engines on TSMC 28nm. [Uncaptioned image]
@@
Acceleration for quantized model inference. There have been several prior works on accelerating quantized neural network inference (Qin et al., 2023; Zeng et al., 2024; Reggiani et al., 2023; Chang et al., 2021; Song et al., 2020; Inci et al., 2023; Keller et al., 2022; Yang et al., 2022; Li et al., 2024a; Shen et al., 2024). Mokey (Zadeh et al., 2022) and Olive (Guo et al., 2023) apply outlier-aware quantization methods to transformer-based LLMs. LUT-GEMM (Park et al., 2024c) proposes a lookup-table-based GPU kernel to eliminate dequantization overhead in quantized LLM inference. Tender (Lee et al., 2024b) quantizes KV cache as well as weights and activations, but its accuracy loss is significant. In contrast, Oaken is an LLM inference acceleration solution optimized for KV cache quantization, offering a scalable solution while minimizing accuracy loss.
@@ 8. Conclusion
Acknowledgements.
We thank the anonymous reviewers for their comments and feedback. This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) (No.2018-0-00503, No.RS-2024-00459797, Development of ML compiler framework for on-device AI), IITP under the Graduate School of Artificial Intelligence Semiconductor (IITP-2025-RS-2023-00256472), and the National Research Foundation of Korea (NRF) (RS-2024-00342148), grant funded by the Korea government (MSIT). This work was also partly supported by HyperAccel.